OpenStudy (anonymous):

find the value of 2log10^100

OpenStudy (anonymous):

the answer is 2

OpenStudy (anonymous):

\(\large 2log10^{100}=2*100*log10=2*100*1=200 \)

OpenStudy (anonymous):

yah the 2 00's didn't send :/ sorry bout that...